Session: May 2002 (session 2)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Carlo Migoni at Thu May 30 00:35:13 CEST 2002) Success

Noto (P. Leto at Thu May 30 13:31:41 CEST 2002) Success
No known problems

Jodrell_Bank (Alastair Gunn at Thu May 30 16:29:16 CEST 2002) Minor failures
A power supply problem has led to VCs loosing lock for some periods of time. VC08 unlocked and swapped with VC09 about 21:12 UT. By 22:00 UT VCs 2,3,4,5 and 6 all loosing lock for substantial periods of time. Data quality unknown. Power supply problem is being investigated.

Torun (Kazimierz Borkowski at Fri May 31 07:14:57 CEST 2002) Success
No known problems.

Effelsberg (A. Kraus at Fri May 31 12:39:55 CEST 2002) Minor failures
We had a problem with the frontend control unit: No cal signal until UT1730.

Onsala (John Conway at Mon Jun 3 14:27:04 CEST 2002) Success
No known problems